Q.1.1 Hybrid Dealers bought computer equipment for R54 000 on 1 July 2018. The equipment is depreciated at 20% according to the diminishing balance method.


Required:

Prepare the depreciation and closing transfer entries in the general journal of

Hybrid Dealers on 31 March 2019 and 31 March 2020. (15)


1
Expert's answer
2021-01-12T07:56:46-0500

Answer:


Step 1:


Depreciation means the loss in value of assets because of usage of assets , passage of time or change in technology.

Since , only 20% is given and not 20% per annum , hence we are charging depreciation at 20% without considering the months.

If 20% per annum is given then depreciation for the year ended 31st March , 2019 will be charged for nine months and next year for 12 months.


Step 2

Cost of Computer equipment                                                                                        = 54,000

Less :- Depreciation for the year 31st March 2019 "("54000*"\dfrac {20} {100}"")"                          = (10,800)          

Written down value as on 1st April , 2019 "("54000-10800")"                                         = 43,200

Less :- Depreciation for the year 31st March , 2020  "("43200*"\dfrac {20} {100}"")"        = (8,640)                            

Written down value as on 1st April ,2020 "("43200-8640")"         = 34,560


Step 3